OpenStudy (anonymous):

Given that the pair of triangles is similar, find the length of the side labeled x. 1st triangle is 2.5 & 1.5 2nd triangle is x & 0.75

OpenStudy (anonymous):

similar to last post 2.5/x = 1.5 / 0.75 2.5 / x = 2 now its easy to find x

OpenStudy (anonymous):

how

OpenStudy (anonymous):

multiply both sides of the equation by x 2.5 * x -- = 2x x 2.5 = 2x divide both sides by 2 1.25 = x

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Ratio rule: x = 2.5 * .75/ 1.5 = 1.25

OpenStudy (anonymous):

Poor you, jimmy :P